During startup of a power plant, a turbine accelerates from rest at 0.54 rad/s2

1- How long does it take to reach its 4000 rpm operating speed?

2- How many revolutions does it make during this time?
Answer #1

here,

accelratiob , alpha = 0.54 rad/s^2

a)

let the time taken be t

operating speed , w = 4000 rpm = 418.67 rad/s

w = 0 + alpha * t

418.67 = 0.54 * t

t = 775.3 s

b)

theta = w^2 /2*alpha

theta = 418.67^2 /( 2 * 0.54)

theta = 1.62 * 10^5 rad

the number of revolutions , n = theta /2pi = 2.58 * 10^4 rev